For a journal entry with only two lines, the following entry is valid:


  1. For a journal entry with only two lines, the following entry is valid: Decrease in Owners' Equity, Increase in Expense.

    1. False                                                                      B) True

  1. Allocation is primarily as issue in

    1. Measuring Income                                                B) None of the other three

C) Measuring Assets                                                  D) Both Measuring Assets and Income

  1. ABC has beginning inventory for the year of $18,000. During the year, ABC purchases inventory for $230,000 and has cost of goods sold equal to $233,000. ABC's ending inventory equals:

A) $21,000.                       B) $18,000.                      C) $15,000.                                          D) $19,000.

  1. Of the following inventories 1) Raw Materials Inventory, 2) Work in Process Inventory, 3) Finished Goods Inventory, how many turn into Cost of Goods Sold directly?

    1. Two                              B) None                            C) One D) Three

  2. ABC buys widgets for $5 cash and sells them on account for $8. What is the sacrifice value of a receivable on the books of ABC?

    1. $8

    2. $5

    3. $3

    4. Impossible to determine from the given information

  3. For a journal entry with only two lines, the following entry is valid: Increase in Revenue, Decrease in Expense.

    1. False                                                                      B) True

  4. In the P/E ratio

    1. P is forward looking; E is backward looking.

    2. P is backward looking; E is backward looking.

    3. P is forward looking; E is forward looking.

    4. P is backward looking; E is forward looking.

  5. Cost of goods sold $420,000   Sales revenue800,000   Nonoperating expenses 10,000  Operating expenses 170,000   Income tax expense 80,000

    370)

    What is operating income?

    A) $200,000.                     B) $210,000.                    C) $380,000.                                          D) $120,000.
Answer #1
1- For a journal entry with only two lines, the following entry is valid: Decrease in Owners' Equity, Increase in Expense FALSE
2- Measuring income
3- cost of goods sold = opening inventory+purchase-closing inventory option is C -15000
233000= 18000+230000-closing inventory closing inventory = 248000-233000= 15000
4- Option is D - 3 all the three turn into cost of goods sold directly
5- Option is B -5
6- FALSE
7- Option is 2 P is backward looking; E is backward looking.
8- Option is B   sales-cost of goods sold- operating expenses 210000